# Advanced Encryption Standard AI artificial intelligence

### v1

@@
Stealing Maggie’s Secrets—On the Challenges of IP Theft
Through FPGA Reverse Engineering
Simon Klix1[Uncaptioned image], Nils Albartus1[Uncaptioned image], Julian Speith1[Uncaptioned image], Paul Staat1[Uncaptioned image], Alice Verstege2[Uncaptioned image], Annika Wilde2[Uncaptioned image],
Daniel Lammers2[Uncaptioned image], Jörn Langheinrich1[Uncaptioned image], Christian Kison3[Uncaptioned image], Sebastian Sester3[Uncaptioned image],
Daniel Holcomb4[Uncaptioned image], Christof Paar12[Uncaptioned image]
1Max Planck Institute for Security and Privacy, Bochum, Germany 2Ruhr University Bochum, Horst Görtz Institute for IT Security, Bochum, Germany 3Bundeskriminalamt, Wiesbaden, Germany 4UMass Amherst, MA, USA
Abstract
@@
Intellectual Property (IP) theft is a cause for major financial and reputational damage, reportedly in the range of hundreds of billions of dollars annually in the U.S. alone. Field Programmable Gate Arrays (FPGAs) are particularly exposed to IP theft, because their configuration file contains the IP in a proprietary format that can be mapped to a gate-level netlist with moderate effort. Despite this threat, the scientific understanding of this issue lacks behind reality, thereby preventing an in-depth assessment of IP theft from FPGAs in academia. We address this discrepancy through a real-world case study on a Lattice iCE40 FPGA found inside iPhone 7. Apple refers to this FPGA as Maggie. By reverse engineering the proprietary signal-processing algorithm implemented on Maggie, we generate novel insights into the actual efforts required to commit FPGA IP theft and the challenges an attacker faces on the way. Informed by our case study, we then introduce generalized netlist reverse engineering techniques that drastically reduce the required manual effort and are applicable across a diverse spectrum of FPGA implementations and architectures. We evaluate these techniques on seven benchmarks that are representative for different FPGA applications and have been synthesized for Xilinx and Lattice FPGAs. Finally, we provide a comprehensive open-source tool-suite of netlist reverse engineering techniques to foster future research, enable the community to perform realistic threat assessments, and facilitate the evaluation of novel countermeasures.
Index Terms:
@@
IP theft, reverse engineering, bitstream, gate-level netlist, FPGA, iPhone
1 Introduction
@@
Many companies invest heavily into creating intellectual property (IP), which often forms the backbone of their economic value and competitive advantage. For instance, Apple alone spent over \$26 billion on research and development (R&D) in 2022 [1]. Not surprisingly, valuable IP can nowadays often be found within electronics and integrated circuits in particular. Here, it can take the form of proprietary algorithms for domains like digital signal processing (DSP), CPUs, or security functions. As a flip side of this development, IP theft, e.g., by competitors or hostile nation states, has become a major issue. In the U.S. alone, IP theft causes hundreds of billions of dollars in damages every year, with the semiconductor market being of particular concern [2]. In this paper, we look at an important aspect of this issue, namely assessing the security of IP in field-programmable gate arrays, which are, as we argue, particularly vulnerable against IP theft.
@@
In many highly specialized domains—such as spaceflight, military communications, network routers, or medical devices—FPGAs are a vital component, often incorporating proprietary IP. FPGAs are (re-)configurable logic devices that are programmed using a bitstream with a proprietary file format. This bitstream is generated by electronic design automation (EDA) tools that are provided by the FPGA vendors themselves. These tools map a high-level design, written in a hardware description language (HDL) such as Verilog, to a bitstream during synthesis. At the end of this process, the implemented IP is fully encapsulated in the bitstream file. This file is then stored externally to the FPGA, but residing on the target device.
@@
To better assess the threat of, and defend against, real-world IP theft from FPGA implementations, it is crucial to improve our understanding of the associated challenges and required efforts. In most cases, hardware IP theft entails reverse engineering the design in order to recover the implemented algorithms and their instantiation.
@@
Modern nanometer-scale ICs come with a natural defense against reverse engineering due to the high cost and labor-intensive nature of the netlist extraction process [3, 4, 5, 6, 7]—a financial barrier that deters all but the most resourceful entities, such as nation-states. However, the bar for FPGA reverse engineering is much lower. To this end, an attacker would need to gain access to the bitstream and map it to a gate-level netlist, a process that is already well understood in research despite the proprietary nature of the bitstream format [8, 9, 10, 11, 12, 13, 14, 15]. FPGA vendors commonly provide cryptographic bitstream protections as a safeguard for valuable IP. Nonetheless, these measures have repeatedly been demonstrated to be susceptible to side-channel attacks [16, 17, 18, 19, 20] or protocol flaws [21, 22]. Interestingly, many low-cost FPGAs do not even offer such protections. Even if these protective measures are available and would withstand such attacks, they must be proactively enabled by the FPGA developer. Hence, we assume access to the plaintext bitstream, and thus the gate-level netlist, is generally achievable in practice.
@@
In this situation, a reverse engineer is typically faced with a gate-level netlist that lacks any hierarchy, module boundaries, and word-level information such as data types, see Section 2.3. Many netlist reverse engineering techniques have tried to address these challenges in recent years [23], but they often deal only with isolated sub-problems such as register recovery, lack reference implementations, and use unrealistic and/or outdated benchmarks, see Section 5.3. In particular, an end-to-end analysis of a real-world FPGA design remains largely unexplored in literature. As a consequence, the actual threat potential of FPGA IP theft as well as the unique challenges it entails remain unknown. Thus, the motivation of the paper at hand is to investigate the reverse engineering of a complex FPGA netlist that is recovered from a real-world device in a black-box setting. The goal is to obtain and analyze the algorithmic description of the design. A better understanding of this issue can aid in the design of sound defenses and inform risk analysis of industry as well as governments. We note that FPGA reverse engineering also paves the way for attacks other than IP theft, e.g., retrieving hard-coded cryptographic keys or inserting hardware Trojans.
@@
For a realistic assessment of the challenges of real-world FPGA reverse engineering in the context of IP theft, we first set out to answer the following research question:
RQ1 What challenges and efforts are entailed with real-world IP theft through FPGA reverse engineering in a black-box setting?
To answer RQ1, we demonstrate the steps and efforts required to extract IP from an FPGA implementation through a case study on a Lattice iCE40 FPGA found within the iPhone 7. Apple refers to this FPGA as Maggie in the iPhone firmware. To this end, we combine existing techniques and propose new approaches to recover an algorithmic description of the target FPGA implementation, effectively extracting the IP controlling the iPhone’s Taptic Engine, see Section 3. In addition to evaluating existing techniques and proposing novel approaches to FPGA reverse engineering, we provide unique insights that foster developing better defenses against IP theft. Based on our findings, a second research question arises naturally: RQ2 To what extent can FPGA reverse engineering be generalized across architectures and implementations?
In response to RQ2, we generalize the methods developed for our case study. This yields automated methods that are applicable across FPGA architectures and implementations. In particular, we target Xilinx 7-Series and Lattice iCE40 FPGA families, see Section 4. We evaluate the presented techniques on a diverse set of benchmark netlists and discuss differences in effectiveness. In line with general efforts of the security community to provide open-source implementations, thereby removing financial barriers and enabling third-party verification, we will share our benchmarks and implementations as plugins to HAL [24] upon publication of this work. Thereby, we hope to encourage more insight into FPGA reverse engineering in follow-up work by academia and industry, and foster the development of countermeasures, see Section 5.
@@
Despite not revealing any security vulnerabilities, we informed Apple and Lattice Semiconductor about our findings prior to publication.
2 Attacker Model & Challenges
We now present our attacker model, introduce relevant terminology, and discuss the challenges of netlist reverse engineering encountered in our case study.
2.1 Attacker Model
Our attacker model revolves around IP theft, which is a common threat considered in hardware security research [25, 26]. An attacker obtains a device featuring an FPGA and aims to recover the contained IP by extracting an algorithmic model of the design implemented on the FPGA by means of reverse engineering. We assume the attacker has gained access to the bitstream configuring the FPGA by either extracting it from the device or from firmware. Given that the attacker possesses the device, they can execute the FPGA implementation and modify the hardware as needed.
@@ 2.1 Attacker Model
2.2 Netlist Reverse Engineering Terminology
@@
2.3 Challenges of Netlist Reverse Engineering
Committing IP theft on an FPGA requires analyzing the gate-level netlist extracted from the bitstream, posing unique challenges. These challenges arise from the need to reconstruct information removed during synthesis. The challenges we faced in our case study include:
  1. C1
@@ 2.3 Challenges of Netlist Reverse Engineering
  3. C3
    Synthesizer Optimizations. Synthesizer optimizations can result in similar operations on HDL level looking vastly different in a gate-level netlist. Combinational logic may be merged across module boundaries and gates or even entire sub-circuits may be duplicated to facilitate a high fan-out. Such optimizations always depend on the available standard cells, timing and area constraints, routing requirements, and varying optimization strategies.
  4. C4
@@
  5. C5
    Data Dependency. The behavior of the control path—and in extension the data path—depended on external data and instructions fed to the FPGA as input. Hence, analyzing such a netlist without input data is challenging, especially if the control path cannot be properly dissected.
  6. C6
@@
  7. C7
    Semantic Analysis. Extracting a word-level algorithmic representation leaves a reverse engineer with the task of assigning meaning and symbols to different values of the computation. To gain a high-level understanding, including the underlying rationale and specific design decisions, the functionality needs to be dissected rigorously from the system-level down to individual building blocks, calling for domain-specific expertise.
3 Case Study on iPhone 7
@@
The iPhone 7 and iPhone 7 Plus were released in September 2016. The smartphones feature a novel capacitive home button with haptic feedback being generated by the Taptic Engine first introduced in Apple Watch and iPhone 6s. Notably, both iPhone 7 and iPhone 7 Plus contain a Lattice iCE5LP4K-SWG36 FPGA. Apple refers to this chip as Maggie throughout firmware and leaked printed circuit board (PCB) schematics. There are many speculations revolving around the purpose of Maggie [27, 28, 29].
Figure 1: Maggie as part of the Taptic Engine controller.
@@
From analyzing leaked PCB schematics, the iPhone firmware, and patents granted to Apple [30, 31], we concluded that Maggie controls the Taptic Engine, and determined its I/O connectivity and communication interfaces, see Figure 1. This alone does not enable effective IP theft, as we still lack implementation details on the instantiated signal processing algorithms and their parameters. Assuming the role of the attacker, we therefore need to investigate the circuit implemented on Maggie to fully recover Apple’s Taptic Engine IP. To this end, we split the iPhone FPGA reverse engineering process into four steps, see Figure 2.
Figure 2: Overview of our case study on iPhone 7.
3.1 Step 1: Netlist Recovery
@@
The bitstream format is typically a secret well-kept by FPGA vendors, although everyone can generate bitstreams using respective EDA tools. Project IceStorm [32] entails a comprehensive documentation database of the Lattice iCE40 bitstream format generated by fuzzing the Lattice iCEcube2 EDA tool. Project IceStorm is part of F4PGA [33], an open-source EDA tool-flow developed as an alternative to proprietary vendor software. As part of our efforts, we extended Project IceStorm by increasing its fuzzing coverage and adding support for the pinout of Maggie’s SWG36 chip package.
@@
Given a database detailing the mapping from bits to netlist elements, an arbitrary bitstream can be converted back to a gate-level netlist in full automation. In light of our case study, we developed a custom bitstream conversion tool that translates any Lattice iCE40 bitstream into a gate-level netlist. Our tool depends on the Project IceStorm database to be complete and correct. For Maggie, we observed 36 nets with missing sources or destinations, hinting at an incomplete bitstream database. Later on, we manually repaired these nets using information on surrounding circuitry revealed once our analysis progressed.
The extracted netlist contains 5046 gates comprising 3241 look-up tables, 422 carry gates, 1342 flip-flops, 18 block-s, and 4 DSP units.
3.2 Step 2: Word-Level Reconstruction
After extracting the gate-level netlist, we now aim to recover word-level structures from the unstructured sea-of-gates using the netlist reverse engineering framework HAL [24]. To this end, we were confronted with the challenges outlined in Section 2.3.

### v2

@@
Stealing Maggie’s Secrets—On the Challenges of IP Theft Through FPGA Reverse Engineering
Simon Klix $0$-0002-9369-2901 MPI-SPBochumGermany Nils Albartus $0$-0003-2449-1134 MPI-SPBochumGermany Julian Speith $0$-0002-8408-8518 MPI-SPBochumGermany Paul Staat $0$-0002-7539-4847 MPI-SPBochumGermany Jörn Langheinrich $0$-0002-8583-5503 MPI-SPBochumGermany Alice Verstege 0009-0005-2807-0806 Ruhr University BochumBochumGermany Annika Wilde $0$-0002-6092-7866 Ruhr University BochumBochumGermany Daniel Lammers $0$-0002-9134-8568 Ruhr University BochumBochumGermany Christian Kison $0$-0002-5830-7692 BundeskriminalamtWiesbadenGermany Sebastian Sester-Wehle $0$-0002-4938-2084 BundeskriminalamtWiesbadenGermany Daniel Holcomb $0$-0002-2052-9820 UMass AmherstAmherstMAUSA  and  Christof Paar $0$-0001-8681-2277 MPI-SPBochumGermany
Abstract.
@@
Intellectual Property (IP) theft is a cause of major financial and reputational damage, reportedly in the range of hundreds of billions of dollars annually in the U.S. alone. Field Programmable Gate Arrays (FPGAs) are particularly exposed to IP theft, because their configuration file contains the IP in a proprietary format that can be mapped to a gate-level netlist with moderate effort. Despite this threat, the scientific understanding of this issue lacks behind reality, thereby preventing an in-depth assessment of IP theft from FPGAs in academia. We address this discrepancy through a real-world case study on a Lattice iCE40 FPGA found inside iPhone 7. Apple refers to this FPGA as Maggie. By reverse engineering the proprietary signal-processing algorithm implemented on Maggie, we generate novel insights into the actual efforts required to commit FPGA IP theft and the challenges an attacker faces on the way. Informed by our case study, we then introduce generalized netlist reverse engineering techniques that drastically reduce the required manual effort and are applicable across a diverse spectrum of FPGA implementations and architectures. We evaluate these techniques on six benchmarks that are representative of different FPGA applications and have been synthesized for Xilinx and Lattice FPGAs, as well as in an end-to-end white-box case study. Finally, we provide a comprehensive open-source tool suite of netlist reverse engineering techniques to foster future research, enable the community to perform realistic threat assessments, and facilitate the evaluation of novel countermeasures.
@@
FPGA reverse engineering, IP theft, bitstream, gate-level netlist
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ACM Conference on Computer and Communications Security; Oct. 14-18, 2024; Salt Lake City, UT, USAisbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacy Hardware reverse engineering
1. Introduction
@@
Many companies invest heavily into creating intellectual property (IP), which often forms the backbone of their economic value and competitive advantage. For instance, Apple alone spent over \$26 billion on research and development (R&D) in 2022 (Insider Monkey and Macrotrends, 2022). Not surprisingly, valuable IP can nowadays often be found within electronics and integrated circuits in particular. It can take the form of proprietary algorithms for domains like digital signal processing (DSP), CPUs, or security functions. As a flip side of this development, IP theft, e.g., by competitors or hostile nation states, has become a major issue. In the U.S. alone, IP theft causes hundreds of billions of dollars in damages every year, with the semiconductor market being of particular concern (Li, 2020). In this paper, we look at an important aspect of this issue, namely assessing the security of IP in field-programmable gate arrays, which are, as we argue, particularly vulnerable against IP theft.
@@
In many highly specialized domains—such as spaceflight, military communications, network routers, or medical devices—FPGAs are a vital component, often incorporating proprietary IP. FPGAs are (re-)configurable logic devices that are programmed using a bitstream with a proprietary file format. This bitstream is generated by electronic design automation (EDA) tools provided by the FPGA vendors. These tools map a high-level design, written in a hardware description language (HDL) such as Verilog, to a bitstream during synthesis. At the end of this process, the implemented IP is fully encapsulated in the bitstream file. This file is stored externally to the FPGA but resides on the target device.
@@
Improving our understanding of the associated challenges and required efforts is crucial to better assess the threat of, and defend against, real-world IP theft from FPGA implementations. In most cases, hardware IP theft entails reverse engineering the design in order to recover the implemented algorithms and their instantiation.
@@
Modern nanometer-scale ICs come with a natural defense against reverse engineering due to the high cost and labor-intensive nature of the netlist extraction process (Lippmann et al., 2019; Fyrbiak et al., 2017; Quadir et al., 2016; Torrance and James, 2011, 2009)—a financial barrier that deters all but the most resourceful entities, such as nation-states. However, the bar for FPGA reverse engineering is much lower. To this end, an attacker would need to access the bitstream and map it to a gate-level netlist. This process is already well understood despite the proprietary nature of the bitstream format (Ender et al., 2019; Ziener et al., 2006; Note and Rannaud, 2008; Benz et al., 2012; Ding et al., 2013; Pham et al., 2017; Note, 2008; Alliance, 2023b). FPGA vendors commonly provide cryptographic bitstream protections as a safeguard for valuable IP. Nonetheless, these measures have repeatedly been demonstrated to be susceptible to side-channel attacks (Moradi et al., 2011, 2012; Moradi and Schneider, 2016; Swierczynski et al., 2015b; Tajik et al., 2017) or protocol flaws (Ender et al., 2020, 2022). Many low-cost FPGAs do not even offer such protections. Even if these protective measures are available and resist such attacks, they must be proactively enabled by the FPGA developer. Hence, we assume access to the plaintext bitstream, and thus the gate-level netlist, is generally achievable in practice.
@@
In this situation, a reverse engineer is typically faced with a gate-level netlist that lacks any hierarchy, module boundaries, and word-level information such as data types, see Section 2.3. Many netlist reverse engineering techniques have tried to address these challenges in recent years (Azriel et al., 2021), but they often deal only with isolated sub-problems such as register recovery, lack reference implementations, and use unrealistic and/or outdated benchmarks, see Section 5.3. In particular, an end-to-end analysis of a real-world FPGA design remains largely unexplored in literature. As a consequence, the actual threat potential of FPGA IP theft as well as the unique challenges it entails remain unknown. Thus, the motivation of the paper at hand is to investigate the reverse engineering of a complex FPGA netlist that is recovered from a real-world device in a black-box setting. The goal is to obtain and analyze the algorithmic description of the design. A better understanding of this issue can aid in the design of sound defenses and inform risk analysis of industry as well as governments. FPGA reverse engineering also paves the way for attacks other than IP theft, e.g., retrieving hard-coded cryptographic keys or inserting hardware Trojans.
@@
For a realistic assessment of real-world FPGA reverse engineering challenges in the context of IP theft, we first set out to answer the following research question:
{mdframed}
  1. RQ1:
    What challenges and efforts are entailed with IP theft through FPGA reverse engineering in a black-box setting?
To answer RQ1, we demonstrate the steps and efforts required to extract IP from an FPGA implementation through a case study on a Lattice iCE40 FPGA found within the iPhone 7. Apple refers to this FPGA as Maggie in the iPhone firmware. To this end, we combine existing techniques and propose new approaches to recover an algorithmic description of the target FPGA implementation, effectively extracting the IP controlling the iPhone’s Taptic Engine, see Section 3. In addition to evaluating existing techniques and proposing novel approaches to FPGA reverse engineering, we provide unique insights that foster developing better defenses against IP theft. Based on our findings, a second research question arises:
{mdframed}
  1. RQ2:
    To what extent can FPGA reverse engineering be generalized across architectures and implementations?
In response to RQ2, we generalize the methods developed for our case study. This yields automated methods that are applicable across FPGA architectures and implementations. In particular, we target Xilinx 7-Series and Lattice iCE40 FPGA families, see Section 4. We evaluate the presented techniques in isolation on a diverse set of benchmark netlists and discuss differences in effectiveness. Additionally, we conduct an end-to-end white-box case study to demonstrate the effectiveness of our workflow as a whole. In line with the efforts of the security community to provide open-source implementations, thereby removing financial barriers and enabling third-party verification, we share our benchmarks and implementations as plugins to HAL (HAL, 2018). Thereby, we hope to encourage more insight into FPGA reverse engineering in follow-up work by academia and industry, and foster the development of countermeasures, see Section 5.
@@ 1. RQ2:
Despite not revealing any security vulnerabilities, we informed Apple and Lattice Semiconductor about our findings before publication. In agreement with the vendors, we will not publish the gate-level netlist recovered from Maggie’s bitstream or our bitstream conversion tooling for Lattice iCE40 FPGAs.
2. Attacker Model & Challenges
2.1. Attacker Model
Our attacker model revolves around IP theft, which is a common threat considered in hardware security research (Forte et al., 2017; Shamsi et al., 2017). An attacker obtains a device featuring an FPGA and aims to recover the contained IP by extracting an algorithmic model of the design implemented on the FPGA by means of reverse engineering. We assume the attacker has gained access to the bitstream configuring the FPGA by either extracting it from the device or from firmware. Given that the attacker possesses the device, they can execute the FPGA implementation and modify the hardware as needed.
@@ 2.1. Attacker Model
2.2. Netlist Reverse Engineering Terminology
@@
2.3. Challenges of Netlist Reverse Engineering
Committing IP theft on an FPGA requires analyzing the gate-level netlist extracted from the bitstream, posing unique challenges. These challenges arise from the need to reconstruct information removed during synthesis. The challenges we faced include:
  1. C1:
@@ 2. C2:
  3. C3:
    Synthesizer Optimizations. Synthesizer optimizations can result in similar operations on HDL level looking vastly different in a gate-level netlist. Logic may be merged across module boundaries and gates or even entire sub-circuits may be duplicated to facilitate a high fan-out. Such optimizations always depend on the available standard cells, timing and area constraints, routing requirements, and varying optimization strategies.
  4. C4:
@@
  5. C5:
    Data Dependency. In our case study, the behavior of the control path—and in extension the data path—depended on external data and instructions fed to the FPGA as input. Hence, analyzing such a netlist without input data is challenging, especially if the control path cannot be properly dissected.
  6. C6:
@@
  7. C7:
    Semantic Analysis. Extracting a word-level algorithmic representation leaves a reverse engineer with the task of assigning meaning and symbols to different values of the computation. To gain a high-level understanding, including the underlying rationale and specific design decisions, the functionality needs to be dissected rigorously from the system level down to individual building blocks, calling for domain-specific expertise.
3. Case Study on iPhone 7
@@
The iPhone 7 and iPhone 7 Plus were released in September 2016. The smartphones feature a novel capacitive home button with haptic feedback being generated by the Taptic Engine first introduced in Apple Watch and iPhone 6s. Notably, both iPhone 7 and iPhone 7 Plus contain a Lattice iCE5LP4K-SWG36 FPGA. Apple refers to this chip as Maggie throughout firmware and leaked printed circuit board (PCB) schematics. There are many speculations revolving around the purpose of Maggie (Martellaro, 2016; Tilley, 2016; iFixit, 2016).
Figure 1. Maggie as part of the Taptic Engine controller.
[…]
[¡short description¿]¡long description¿
@@
From analyzing leaked PCB schematics, the iPhone firmware, and patents granted to Apple (Hajati, 2021; Hajati and Patel, 2018), we concluded that Maggie controls the Taptic Engine, and determined its I/O connectivity and communication interfaces, see Figure 1. This alone does not enable effective IP theft, as we still lack implementation details on the instantiated signal processing algorithms and their parameters. Assuming the role of the attacker, we therefore need to investigate the circuit implemented on Maggie to fully recover Apple’s Taptic Engine IP. To this end, we split the iPhone FPGA reverse engineering process into four steps, see Figure 2.
Figure 2. Overview of our case study on iPhone 7.
3.1. Step 1: Netlist Recovery
@@
The bitstream format is typically a secret well-kept by FPGA vendors, although everyone can generate bitstreams using respective EDA tools. Project IceStorm (Alliance, 2023a) entails a comprehensive documentation database of the Lattice iCE40 bitstream format generated by fuzzing the Lattice iCEcube2 EDA tool. Project IceStorm is part of F4PGA (Alliance, 2022), an open-source EDA tool-flow developed as an alternative to proprietary vendor software. As part of our efforts, we extended Project IceStorm by increasing its fuzzing coverage and adding support for the pinout of Maggie’s SWG36 chip package.
@@
Given a database detailing the mapping from bits to netlist elements, an arbitrary bitstream can be converted back to a gate-level netlist in full automation. In light of our case study, we developed a custom bitstream conversion tool that translates any Lattice iCE40 bitstream into a gate-level netlist. Our tool depends on the Project IceStorm database to be complete and correct. For Maggie, we observed 36 nets with missing sources or destinations, hinting at an incomplete bitstream database. Later on, we manually repaired these nets using information on surrounding circuitry revealed once our analysis progressed. The extracted netlist contains 5046 gates comprising 3241 look-up tables, 422 carry gates, 1342 flip-flops, 18 block-s, and 4 DSP units.
3.2. Step 2: Word-Level Reconstruction
After extracting the gate-level netlist, we aimed to recover word-level structures from the unstructured sea-of-gates using the netlist reverse engineering framework HAL (HAL, 2018). To this end, we were confronted with the challenges outlined in Section 2.3.